The lenght of a string between a kite and Point A on the ground is 442meter. If the height of the kite is 42metwr from the ground.Find the horizontal angle and distance of the kite from point A.

1
Expert's answer
2022-02-13T12:15:09-0500

"\alpha=\arctan \frac {\sqrt{b^2-a^2}}a=84.5°,"

"c=\sqrt{b^2-a^2}=440~m."